The piston of a hydraulic automobile lift is 30 cm in diameter. What pressure, in

pascals, is required to lift a car of mass 1500 kg. Also express this pressure in

atmosphere.


1
Expert's answer
2022-03-06T15:18:36-0500

The pressure

"P=\frac{F}{A}=\frac{mg}{\pi d^2/4}"

"P=\frac{1500*9.8}{3.14*0.3^2/4}=2.08*10^5\:\rm Pa=2.05\: atm"